5. Calculating tax incidence

Suppose that the U.S. government decides to charge cola producers a tax. Before the tax, 30 billion cases of cola were sold every year at a price of $5 per case. After the tax, 24 billion cases of cola are sold every year; consumers pay $6 per case, and producers receive $2 per case (after paying the tax).

The amount of the tax on a case of cola is _____per case. Of this amount, the burden that falls on consumers is______per case, and the burden that falls on producers is_____per case.

True or False: The effect of the tax on the quantity sold would have been the same as if the tax had been levied on consumers.

True

False
Answer #1

1) tax = price paid by consumers - price received by producers

= 6-2

= $ 4

2) tax burden on consumers = (6-5) = $ 1

3) burden on producers = (5-2) = $ 3 per case

4) true

The effect of tax will be same, whether it is applied on consumers or producers